OpenStudy (anonymous):

a long distance trucker traveled 180 miles in one direction during a snow storm. the return trip in rainy weather was accomplished at double the speed and took 3 hours less time. find the speed going.

OpenStudy (anonymous):

a t =180 (2a)*(t-3)=180 solve for a

OpenStudy (anonymous):

did you find a?

OpenStudy (anonymous):

no I need help

OpenStudy (anonymous):

(2a)*(t-3)=180 2at-6a=180 we know at=180 ,so 2 at is 360 360-6a=180 -6a=-180 a=30